What is the probability that a student who is guessing at answers gets the first three questions on a true-false test correct?
Answer by stanbon(57984)

What is the probability that a student who is guessing at answers gets the first three questions on a true-false test correct?
-------------------
Prob of getting each one correct = 1/2
----------------
P(getting 3 correct) = (1/2)^3 = 1/8
======================================
Cheers,
Stan H.